Engineering Night UT 20180331

Night Plan:

  • Before Opening:
    • DONE PCS Patch - IT 7084 (Petr)
      • No further on-sky test of this PCS patch is needed beyond confirmation of our ability to track stars.
      • The patch should affect pointing, so if pointing is worse than before or a problem, Petr asked to be called. (OPK)
    • DONE Network Core-A work - IT 7099 (Stephen H. daytime, on-call for night in case of problems)
      • RECOVERY INFO (provided by Matthieu & Stephen H.)
        • The easiest way to recover would simply be to power off core-a (three switches on the back), unless we move the fiber from the singly run switches to it (I would not recommend doing that yet).
        • If we moved the fiber, then it would have to be moved back before power outage.
        • Instrument techs (Elliott and John), have both helped me move fiber from afar before, and I am confident they could also be trusted to do the job.

  • After Opening:
    • DONE AO Startup with OSA (Doug, Al, Steve)
    • DONE Verification of Conditions (Doug, Al, Steve)
    • DONE Total Time for AO startup activities ~ 1 hour

  • IF CLEAR & Seeing < 1.2":
    • AO Engineering Tests - HIGHEST PRIORITY
      • DONE Tip-Tilt Lookup-Table Tests ~4 hours (Doug) - No Arcetri assistance - Plan for Start of Night
      • w-unit transforms for AGW1 and AGW2 ~ confirmation test ~30min (Doug, Barry)
      • W-unit transforms for AGW1 and AGW2 ~ confirmation test ~30 min (Doug, Barry)
    • Field Aberration data for both sides, although DX is the primary target (John H., Barry)
      • The new pointing and collimation models from DougS should be installed prior to collecting field aberration data.

  • ANY Seeing Conditions:
    • DONE PEPSIPFU - final check for IT 6982 ~30 minutes (John H.)
      1. DONE Locate the pointing offsets on a suitable M5 star ~60 deg EL.
      2. DONE Use an ACTIVE/ACTIVE preset to collimate and collect pupil images.
      3. DONE Adjust M3 Tip-Tilt as needed (probably not needed).
      4. DONE Check copointing optics range at ~85 deg EL.
  • Non-Sidereal test - IT 7081 ~ 45 minutes (including overheads) - MODS-Binocular (Prepared by Al & Olga, executed by Jenny and/or Barry)
      • Before 10 UT: 617 Patroclus
      • After 10 UT: JVII Elara

  • Poor Seeing and/or significant clouds (extinction ~ few mags)
    • Collimation and Pointing models -MODS-Binocular-Blue ~ 1-2 hours HIGHEST PRIORITY FOR POOR CONDITIONS
      • The new pointing and collimation models from DougS should be installed prior to collecting more pointing data.
    • TMS Testing - LBC (J. Hill, A. Rakich) ~ any amount of time
    • WRS Testing - LBC (O. Kuhn) ~ any amount of time
    • Poor weather backup programs if TMS and Collimation/Pointing are completed (UT April 1?)


Night

OSA: S. Allanson ISA:

B. Rothberg ISp:

E. Solheid

UT 00:54 - LUCI test darks taken, MODS-1 and MODS-2 awake

UT 01:00 - sievesnap blue only taken

UT 01:15 - preparing for AO startup

UT 01:29 - Enclosure open

Twilight AO Testing

UT 02:48 - AO startup work continuing

UT 03:12 - finished startup testing with LUCI-2

UT 03:55 - finished startup AO testing with both LUCI-s and OSA, setting up for TT LUT work

DX AO TT LUT

UT 04:10 - checking N30fieldstop alignment both LUCIs

UT 04:20 - N30Fieldstop aligned, LUCI-1 off significantly in Y ~ few hundred pixels (X was OK), LUCI-2 ~20-30 pixels offset in X and Y

UT 04:35 - seeing has blown up reading 2.4-3.5" on DIMM

UT 04:41 - Doug has given up in utter despair. Switching to PEPSI for IT 6982 tests, will monitor seeing - if stays poor, move to NS test

PEPSIPFU tests for IT 6982

UT 04:46 Restart oacserver so we can connect to AGw 7 and 8.

Spiralling on Chi UMa at EL=68.

UT 05:08 Stop and Restart OSS

UT 05:20 Now we've got loads (50 arcsec) of copointing range at EL=80. Both pupils are fully on the lenslet arrays.

Change DX hotspot from 341,370 to 361,370. Doug updated DX pupil position.

And we did some guiding for Klaus in horribly variable seeing (0.7 to 2 arcsec FWHM) while we were doing this.

UT 05:41 DX hotspot updated to 355,375.

UT 05:45 Back to AO testing.....

LUCI-AO TT LUT tests for

UT 05:46 - reconfiguring for LUCI-Bino

UT 06:06 - odd shapes on AO WFS (elongated lines)

UT 06:22 - re-aligning N30Fieldstop pinholes

UT 06:32 - Loops closed. Taking LUCI test exposures

UT 06:37 - Run AO completed

UT 06:46 - after tests. 5sec exposures seem best (filter K) - passthru works for NDITS, going to 4 NDITS x 2.51 DIT

UT 06:56 - loops open, at 85 elevation, rotators are truckin'

UT 07:00 - realigning holes, got an error in FM4 on LUCI-1, ran init

UT 07:11 - N30 shifted half an array in readout, changing camera from N30 to N375 and back

UT 07:18 - motors in compensation mirror don't seem to moving anything, turned it off and on, trying again

UT 07:26 - Taking data with LUCI-2 while LUCI-1 FM4 is being troubleshooted

### Instrument: LUCI1

Scale dXi and dYi to N180 pixels:
dXi dYi
Original imaging pixel offsets: -219.00 -20.00 (N375)
Offsets scaled to N180 pixels: -103.22 -9.43

For these N180 camera pixel offsets:

dXimg dYimg dXpup dYpup
-103.22 -9.43 0.00 0.00

Motor offsets to align both the Pupil and FieldStop are:

FM1m1 FM1m2 FM4m1 FM4m2
-846 -7083 -1818 29395
Switched to N30 - the move got us about halfway there, did the AFC align tab, in log FM4 motor 2 moved 12655

Then did a few more minor moves to align N30Fieldstop on LUCI-1

UT 07:51 - LUCI-1 aligned ready for Doug to use

UT 08:48 - moving AO 520 star, slightly fainter to continue AO data collection

UT 09:03 - aligned N30Fieldstop

UT 10:21 - at Elev >82 degrees, having some issues with high elevation and SX side

UT 10:45 - will finish this batch then move to a third set that will have an off-axis AO Ref star - should take us into twilight

UT 10:51 - putting MODS to sleep

UT 11:04 - Moving to 3rd field

UT 11:30 - LUCI-2 N30 fieldstop aligned, will proceed with only LUCI-2 data, blind blind put in LUCI-1

UT 11:43 - SX side loop closed, will start taking LUCI-1 data, K put back

UT 11:50 - issues with opacity? difference of ~1mag between SX and DX, selecting a new target

UT 12:41 - running out of flux

UT 12:51 - Ending AO tests, putting LUCIs in safe mode


-- %USERSIG{BarryRothberg - 2018-03-30}%
Topic revision: r10 - 31 Mar 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback